Find the probability of rolling a 2 on a pair of dice.
1
Expert's answer
2013-02-14T08:53:13-0500
It is only one possible way of rolling a 2:
2=1+1
The probability rolling 1 on one dice:
P(1) = 1/6,
because dice has 6 faces (1,2,3,4,5,6).
Therefore, the probability of rolling a 2 on a pair of dice equal:
P(1)*P(1)= 1/6 * 1/6 = 1/36 = 2.78%
Answer: the probability of rolling a 2 on a pair of dice is 1/36.